A new sales gimmick has 30% of the M&M’s covered with speckles. These “groovy” candies are mixed randomly with normal candies as they are put into the bags for distribution and sale. What is the probability that you will get 2 or fewer speckled Skittles out of 10?

Group of answer choices

38.3%

68.3%

58.3%

48.3%
Answer #1

solution

Given that ,

p = 0.30

1 - p = 1 - 0.30 = 0.70

n = 10

Using binomial probability formula ,

P(X = x) = (n C x) * px * (1 - p)n - x

P(X < 2 ) = P(X=0) +P(X=1)+P(X=2)

= (10 C 0) * 0.300 * (0.70)10+ (10 C 1) * 0.301 * (0.70)9+ (10 C 2) * 0.302 * (0.70)8

answer =38.3%